# Faithful global convergence for the rescaled Consensus–Based Optimization

### v1

@@ 3.1. Single–minimizer
Assume $f(\cdot)$ satisfies Assumptions 2.1 and 3.1, and let $\rho_{*}^{\alpha}$ be the unique invariant probability measure of the rescaled CBO (1.4). Define the probability measure
@@
Let ${\overline{X}}_{.}$ be the solution to the rescaled CBO (1.4). Then in the situation of Theorem 3.3, the expectation of ${\overline{X}}_{.}$ satisfies
$\lim_{t\to\infty}\;{\mathbb{E}}\left[\,{\overline{X}}_{t}\,\right]=\int_{{\mathbb{R}}^{d}}x\rho_{*}^{\alpha}(\mbox{\rm d}x)=\kappa\,\mathfrak{m}_{\alpha}(\rho_{*}^{\alpha})$
and
$\lim_{\alpha\to\infty}\lim_{t\to\infty}\;{\mathbb{E}}\left[\,{\overline{X}}_{t}\,\right]=\kappa\,\lim_{\alpha\to\infty}\mathfrak{m}_{\alpha}(\rho_{*}^{\alpha})=\kappa\,x_{*}$
where $\rho_{*}^{\alpha}$ is the unique invariant probability measure of ${\overline{X}}_{.}$ and $x_{*}$ is the unique global minimizer of $f(\cdot)$.
@@
Let us now consider the (stationary) process (3.4) with initial distribution ${\overline{X}}_{0}\sim\rho_{*}^{\alpha}$. It is well-posed thanks to (2.9). Then taking the expectation in (3.4) and bearing in mind the invariance of $\rho_{*}^{\alpha}$, one has
@@
which together with the previous equality proves the first statement of the corollary.
@@
which holds for any arbitrarily chosen $\varepsilon>0$. In particular, when letting $\varepsilon\to 0$, we obtain the desired result
$\mathfrak{m}_{\alpha}(\rho_{*}^{\alpha})\to x_{*}\quad\mbox{ as }\quad\alpha\to\infty.$
@@ 3.2. Multiple–minimizer
On each $\mathcal{O}_{i}$, the function $f(\cdot)$ is constant equal to its minimal value. When $\mathcal{O}_{i}$ is not a singleton, the graph (landscape) of $f(\cdot)$ is flat, which why these are sometimes called flat minimizers. It is also worth mentioning that a particular example of Assumption 3.5 is when $f(\cdot)$ has $m$ global minimizers $\{x_{*}^{1},\dots,x_{*}^{m}\}=:\mathcal{M}$. In this case, $\mathcal{O}_{i}=\{x_{*}^{i}\}$.
@@
and $\mathcal{M}^{c}_{\varepsilon}={\mathbb{R}}^{d}\setminus\mathcal{M}_{\varepsilon}$ its complement. Here, $\mathrm{dist}(x,\mathcal{O})$ is the Euclidean distance between a point $x\in{\mathbb{R}}^{d}$ and a compact set $\mathcal{O}$, and which is equal to 00 when $x\in\mathcal{O}$.
@@
Property (3.6), which in the case of a single minimizer is (3.1), has already been discussed in [bardi2023eikonal, equation (H)], see also [bardi2023eikonal, Remark 2] for a counter-example.
The first result we need in the sequel is the following tightness of the weighted probability measures.
@@
The next theorem can be viewed as a Laplace principle.
@@
Let $f(\cdot)$ satisfy Assumptions 2.1 and 3.5, and let $\rho_{*}^{\alpha}$ be the unique invariant probability measure of the rescaled CBO (1.4). Let $\eta_{*}^{\alpha}$ be defined as in (3.2). Then, when $\alpha\to\infty$, any weak limit $\eta_{*}$ of the sequence $\{\eta_{*}^{\alpha}\}_{\alpha\geq 1}$ concentrates on $\mathcal{M}$.
@@
As a consequence of the previous theorem, we obtain a convergence in mean, hence also in probability, to the set of minimizers.
@@
Let $Y_{*}^{\alpha}$ be a random variable with probability distribution $\eta_{*}^{\alpha}$ defined in (3.2). In the situation of Theorem 3.9, the following holds
$\lim\limits_{\alpha\to\infty}\,\mathbb{E}\big{[}\mathrm{dist}(Y_{*}^{\alpha},\mathcal{M})\big{]}=0$
where $\mathrm{dist}(Y_{*}^{\alpha},\mathcal{M})=\min\limits_{z\in\mathcal{M}}|Y_{*}^{\alpha}-z|$. In particular, it holds
$\lim\limits_{\alpha\to\infty}\,\mathbb{P}\big{(}\mathrm{dist}(Y_{*}^{\alpha},\mathcal{M})=0\big{)}=1,$
@@
Let $\delta>0$ be fixed. Markov’s inequality yields
$\mathbb{P}\big{(}\mathrm{dist}(Y_{*}^{\alpha},\mathcal{M})>\delta\big{)}\leq\frac{\mathbb{E}\big{[}\mathrm{dist}(Y_{*}^{\alpha},\mathcal{M})\big{]}}{\delta}.$
@@
Then it suffices to show that $\lim\limits_{\alpha\to\infty}\,\mathbb{E}\big{[}\mathrm{dist}(Y_{*}^{\alpha},\mathcal{M})\big{]}=0$.
Using the definition of $Y_{*}^{\alpha}$, we have
@@
$\mathbb{E}\big{[}\mathrm{dist}(Y_{*}^{\alpha},\mathcal{M})\big{]}$ $=\int_{{\mathbb{R}}^{d}}\min\limits_{z\in\mathcal{M}}|y-z|\,\eta_{*}^{\alpha}(\mbox{\rm d}y).$
We claim the following:
(3.9) $\int_{{\mathbb{R}}^{d}}\min\limits_{z\in\mathcal{M}}|y-z|\,\eta_{*}^{\alpha}(\mbox{\rm d}y)\leq\min_{\Delta(\cdot)\in L^{\infty}({\mathbb{R}}^{d},\mathcal{M})}\int_{{\mathbb{R}}^{d}}|y-\Delta(y)|\,\eta_{*}^{\alpha}(\mbox{\rm d}y)$
where $L^{\infty}({\mathbb{R}}^{d},\mathcal{M})$ is the space of measurable functions from ${\mathbb{R}}^{d}$ with values in $\mathcal{M}$.
Therefore, we have
(3.10) $\mathbb{E}\big{[}\mathrm{dist}(Y_{*}^{\alpha},\mathcal{M})\big{]}\leq\int_{{\mathbb{R}}^{d}}|y-\Delta(y)|\,\eta_{*}^{\alpha}(\mbox{\rm d}y)\quad\forall\,\Delta(\cdot)\in L^{\infty}({\mathbb{R}}^{d},\mathcal{M}).$
We can now choose a particular function $\underline{\Delta}(\cdot)$ in the latter inequality, for example
(3.11) $\underline{\Delta}:{\mathbb{R}}^{d}\to\mathcal{M},\quad\quad\underline{\Delta}(y)=\left\{\begin{aligned} &\;\operatorname*{argmin}\limits_{z\in\mathcal{M}}|y-z|,\quad\text{if }y\in\mathcal{M}_{\varepsilon},\\ &\quad z_{*},\quad\text{if }y\in({\mathbb{R}}^{d}\setminus\mathcal{M}_{\varepsilon})=\mathcal{M}_{\varepsilon}^{c},\end{aligned}\right.$
where $z_{*}$ is some point within $\mathcal{M}$ (whose choice is without importance). Recall $\mathcal{M}_{\varepsilon}$ is as defined in Remark 3.6. Such function satisfies $|\underline{\Delta}(y)|\leq\varepsilon+|z_{*}|$. Hence we obtain
$\mathbb{E}\big{[}\mathrm{dist}(Y_{*}^{\alpha},\mathcal{M})\big{]}\leq\int_{{\mathbb{R}}^{d}}|y-\underline{\Delta}(y)|\,\eta_{*}^{\alpha}(\mbox{\rm d}y).$
We split the integral into two terms accordingly
(3.12) $\int_{{\mathbb{R}}^{d}}|y-\underline{\Delta}(y)|\,\eta_{*}^{\alpha}(\mbox{\rm d}y)=\int_{\mathcal{M}_{\varepsilon}}|y-\underline{\Delta}(y)|\,\eta_{*}^{\alpha}(\mbox{\rm d}y)+\int_{\mathcal{M}_{\varepsilon}^{c}}|y-\underline{\Delta}(y)|\,\eta_{*}^{\alpha}(\mbox{\rm d}y).$
@@
The integrand in the first integral is upper-bounded by $\varepsilon$ since $y\in\mathcal{M}_{\varepsilon}$, which means $|y-\underline{\Delta}(y)|\leq\varepsilon$. Thus we have
$\int_{\mathcal{M}_{\varepsilon}}|y-\underline{\Delta}(y)|\,\eta_{*}^{\alpha}(\mbox{\rm d}y)\leq\varepsilon.$
@@
For the second integral in (3.12), we shall proceed as in the second part of the proof of Corollary 3.4. Here, we are working with $y$ in $\mathcal{M}_{\varepsilon}^{c}$, thus $\underline{\Delta}(y)=z_{*}$. Using Cauchy-Schwarz inequality, we have
@@
(3.13) $\int_{\mathcal{M}_{\varepsilon}^{c}}|y-\underline{\Delta}(y)|\,\eta_{*}^{\alpha}(\mbox{\rm d}y)$ $\leq\bigg{(}\eta_{*}^{\alpha}(\mathcal{M}_{\varepsilon}^{c})\bigg{)}^{\frac{1}{2}}\bigg{(}\int_{\mathcal{M}_{\varepsilon}^{c}}|y-\underline{\Delta}(y)|^{2}\,\eta_{*}^{\alpha}(\mbox{\rm d}y)\bigg{)}^{\frac{1}{2}}$
@@
$\leq\bigg{(}\eta_{*}^{\alpha}(\mathcal{M}_{\varepsilon}^{c})\bigg{)}^{\frac{1}{2}}\bigg{(}2\,\int_{\mathcal{M}_{\varepsilon}^{c}}|y|^{2}+|\underline{\Delta}(y)|^{2}\,\eta_{*}^{\alpha}(\mbox{\rm d}y)\bigg{)}^{\frac{1}{2}}$
$\leq\bigg{(}\eta_{*}^{\alpha}(\mathcal{M}_{\varepsilon}^{c})\bigg{)}^{\frac{1}{2}}\bigg{(}2\,|z_{*}|+2\,\eta_{*}^{\alpha}(|\cdot|^{2})\bigg{)}^{\frac{1}{2}}.$
@@
$\mathbb{E}\big{[}\mathrm{dist}(Y_{*}^{\alpha},\mathcal{M})\big{]}\leq\varepsilon+\bigg{(}\eta_{*}^{\alpha}(\mathcal{M}_{\varepsilon}^{c})\bigg{)}^{\frac{1}{2}}\bigg{(}2\,|z_{*}|+2\,\eta_{*}^{\alpha}(|\cdot|^{2})\bigg{)}^{\frac{1}{2}}$
Recalling (2.1) and (2.7), the second moment $\eta_{*}^{\alpha}(|\cdot|^{2})$ is bounded as in (3.5).
Finally, using Theorem 3.9 (see in particular (3.8)), we know that $\eta_{*}^{\alpha}(\mathcal{M}_{\varepsilon}^{c})\,\to 0\,$ as $\alpha\to\infty$, hence $\lim\limits_{\alpha\to\infty}\mathbb{E}\big{[}\mathrm{dist}(Y_{*}^{\alpha},\mathcal{M})\big{]}\leq\varepsilon$. Here $\varepsilon$ being arbitrarily chosen, this yields $\lim\limits_{\alpha\to\infty}\mathbb{E}\big{[}\mathrm{dist}(Y_{*}^{\alpha},\mathcal{M})\big{]}=0$, consequently also $\lim\limits_{\alpha\to 0}\mathbb{P}\big{(}\mathrm{dist}(Y_{*}^{\alpha},\mathcal{M})>\delta\big{)}=0$. The conclusion follows, recalling $\delta$ being arbitrarily chosen.
Proof of the claim (3.9):
For any $\theta>0$, there exists $\Delta_{\theta}(\cdot)\in L^{\infty}({\mathbb{R}}^{d},\mathcal{M})$ such that
$\int_{{\mathbb{R}}^{d}}|y-\Delta_{\theta}(y)|\,\eta_{*}^{\alpha}(\mbox{\rm d}y)\leq\theta+\min_{\Delta(\cdot)\in L^{\infty}({\mathbb{R}}^{d},\mathcal{M})}\int_{{\mathbb{R}}^{d}}|y-\Delta(y)|\,\eta_{*}^{\alpha}(\mbox{\rm d}y).$
We have $\min\limits_{z\in\mathcal{M}}|y-z|\leq|y-\Delta_{\theta}(y)|$. Integrating with respect to $\eta_{*}^{\alpha}$ yields
$\int_{{\mathbb{R}}^{d}}\min\limits_{z\in\mathcal{M}}|y-z|\,\eta_{*}^{\alpha}(\mbox{\rm d}y)\leq\theta+\min_{\Delta(\cdot)\in L^{\infty}({\mathbb{R}}^{d},\mathcal{M})}\int_{{\mathbb{R}}^{d}}|y-\Delta(y)|\,\eta_{*}^{\alpha}(\mbox{\rm d}y).$
This holds for any arbitrarily chosen $\theta$, in particular when $\theta\to 0$, which then proves the claim and concludes the proof. ∎

### v2

@@ 4.1. Single–minimizer
The following result is our first Laplace’s principle in the case of a unique minimizer.
@@
Assume $f(\cdot)$ satisfies Assumptions 2.1 and 4.1, and let $\rho_{*}^{\alpha}$ be the unique invariant probability measure of the rescaled CBO (1.6). Define the probability measure
@@
Assume $f(\cdot)$ satisfies Assumption 2.1. Let $\overline{X}$$_{.}$ be the solution to the rescaled CBO (1.6), and let $\rho_{*}^{\alpha}$ be its unique invariant probability measure. Then
$\lim_{t\to\infty}\;{\mathbb{E}}\left[\,{\overline{X}}_{t}\,\right]=\int_{{\mathbb{R}}^{d}}x\rho_{*}^{\alpha}(\mbox{\rm d}x)=\kappa\,\mathfrak{m}_{\alpha}(\rho_{*}^{\alpha})=\kappa\,\lim\limits_{t\to\infty}\mathfrak{m}_{\alpha}(\rho^{\alpha}_{t}).$
If moreover $f(\cdot)$ satisfies Assumption 4.1, then
$\lim_{\alpha\to\infty}\lim_{t\to\infty}\;\kappa^{-1}\,{\mathbb{E}}\left[\,{\overline{X}}_{t}\,\right]=\lim_{\alpha\to\infty}\lim\limits_{t\to\infty}\mathfrak{m}_{\alpha}(\rho^{\alpha}_{t})=\,\lim_{\alpha\to\infty}\mathfrak{m}_{\alpha}(\rho_{*}^{\alpha})=\,x_{*}.$
@@
Let us first recall that existence and uniqueness of the invariant measure $\rho^{\alpha}_{*}$ are given by Theorem 3.1 and Theorem 3.3 (see also Remark 3.6).
@@
Let us now consider the (stationary) process (4.4) with initial distribution ${\overline{X}}_{0}\sim\rho_{*}^{\alpha}$. It is well-posed thanks to (3.9). Then taking the expectation [^2] Note that this is why the value of $\delta>0$ does not influence the rest of the proof. in (4.4) and bearing in mind the invariance of $\rho_{*}^{\alpha}$, one has
@@
which together with the previous equality yields $\lim_{t\to\infty}\;{\mathbb{E}}\left[\,{\overline{X}}_{t}\,\right]=\kappa\,\mathfrak{m}_{\alpha}(\rho_{*}^{\alpha})$. Finally, using the conclusion of Corollary 3.5, one obtains the first statement.
@@
which holds for any arbitrarily chosen $\varepsilon>0$. In particular, when letting $\varepsilon\to 0$, we obtain the desired result: $\mathfrak{m}_{\alpha}(\rho_{*}^{\alpha})\to x_{*}$ when $\alpha\to\infty.$
@@ 4.2. Multiple–minimizer
On each $\mathcal{O}_{i}$, the function $f(\cdot)$ is constant equal to its minimal value. When $\mathcal{O}_{i}$ is not a singleton, the graph (landscape) of $f(\cdot)$ is flat, which is why these are sometimes called flat minimizers. It is also worth mentioning that a particular example of Assumption 4.5 is when $f(\cdot)$ has $m$ global minimizers $\{x_{*}^{1},\dots,x_{*}^{m}\}=:\mathcal{M}$. In this case, $\mathcal{O}_{i}=\{x_{*}^{i}\}$.
@@
and $\mathcal{M}^{c}_{\varepsilon}={\mathbb{R}}^{d}\setminus\mathcal{M}_{\varepsilon}$ its complement. Here, $\mathrm{dist}(x,\mathcal{O})$ is the Euclidean distance between a point $x\in{\mathbb{R}}^{d}$ and a compact set $\mathcal{O}$, and which is equal to 0 when $x\in\mathcal{O}$.
@@
Property (4.6), which in the case of a single minimizer is (4.1), has already been discussed in [1, equation (H)]; see [1, Remark 2] for a counter-example, and a variation in [2, Assumption (2.3)]. A qualitative interpretation can be found in [39, Assumption (A3)] and it is also used in [38, (A3)] where several examples of optimization problems satisfying this condition are provided. All of the works cited in this remark invoke this property in their respective treatments of global optimization, thereby emphasizing its natural place as a standard assumption.
The first result we need in the sequel is the following tightness of the weighted probability measures. Recall that existence and uniqueness of the invariant measure $\rho_{*}^{\alpha}$ were the object of Theorem 3.1 and Theorem 3.3 (see also Remark 3.6).
@@
The next theorem can be viewed as Laplace’s principle in the case of multiple non-trivial minimizers.
@@
Let $f(\cdot)$ satisfy Assumptions 2.1 and 4.5, and let $\rho_{*}^{\alpha}$ be the unique invariant probability measure of the rescaled CBO (1.6). Let $\eta_{*}^{\alpha}$ be defined as in (4.2). Then for all $\varepsilon>0$, we have
$\eta_{*}^{\alpha}(\mathcal{M}_{\varepsilon})\;\to 1\;\mbox{ when }\;\alpha\to\infty.$
Moreover, when $\alpha\to\infty$, any weak limit $\eta_{*}$ of the sequence $\{\eta_{*}^{\alpha}\}_{\alpha\geq 1}$ concentrates on $\mathcal{M}$.
@@
As a consequence of the previous theorem, we obtain the following convergence when the objective function $f(\cdot)$ has multiple non-trivial minimizers. It indicates that when $f(\cdot)$ admits multiple global minimizers, the consensus point $\mathfrak{m}_{\alpha}(\rho^{\alpha}_{t})$ (or the rescaled mean $\kappa^{-1}\,{\mathbb{E}}\left[\,{\overline{X}}_{t}\,\right]$) of the CBO dynamics converges to a point in $\mathcal{M}$, that is to one of the global minimizers.
@@
In the situation of Theorem 4.9, the following hold
$\lim\limits_{\alpha\to\infty}\,\lim\limits_{t\to\infty}\operatorname{dist}\!\big(\mathfrak{m}_{\alpha}(\rho^{\alpha}_{t}),\,\mathcal{M}\big)=\lim\limits_{\alpha\to\infty}\operatorname{dist}\!\big(\mathfrak{m}_{\alpha}(\rho^{\alpha}_{*}),\,\mathcal{M}\big)=0,$
and $\lim\limits_{\alpha\to\infty}\,\lim\limits_{t\to\infty}\operatorname{dist}\!\left(\kappa^{-1}\,{\mathbb{E}}\left[\,{\overline{X}}_{t}\,\right],\,\mathcal{M}\right)=0.$
@@
Let $\overline{X}$$_{.}$ be the solution to the rescaled CBO (1.6), and let $\rho_{*}^{\alpha}$ be its unique invariant probability measure whose existence and uniqueness are given by Theorem 3.1 and Theorem 3.3 (see also Remark 3.6 for a discussion on the parameters). Besides Assumption 2.1, we are assuming $f(\cdot)$ to satisfy Assumption 4.5 as well, and recall $\mathcal{M}$ being the set of its minimizers. We proceed in several steps.
@@
Step 1. (We show $\lim\limits_{\alpha\to\infty}\operatorname{dist}(\mathfrak{m}_{\alpha}(\rho^{\alpha}_{*}),\,\mathcal{M})=0$.)
Recall
$\eta_{*}^{\alpha}(A)=\frac{\int_{A}e^{-\alpha f(x)}\rho_{*}^{\alpha}(\mbox{\rm d}x)}{\int_{{\mathbb{R}}^{d}}e^{-\alpha f(y)}\rho_{*}^{\alpha}(\mbox{\rm d}y)},\quad A\in\mathcal{B}({\mathbb{R}}^{d})\quad\text{ and }\quad\mathfrak{m}_{\alpha}(\rho^{\alpha}_{*})=\int_{\mathbb{R}^{d}}\,x\;\eta_{*}^{\alpha}(\mbox{\rm d}x).$
We have
(4.9) $\operatorname{dist}(\mathfrak{m}_{\alpha}(\rho^{\alpha}_{*}),\,\mathcal{M})$ $=\inf\limits_{z\in\mathcal{M}}|\mathfrak{m}_{\alpha}(\rho^{\alpha}_{*})-z|$
$=\inf\limits_{z\in\mathcal{M}}\left|\int_{\mathbb{R^{d}}}x\;\eta_{*}^{\alpha}(\mbox{\rm d}x)-z\right|\leq\inf\limits_{z\in\mathcal{M}}\int_{\mathbb{R}^{d}}|x-z|\;\eta_{*}^{\alpha}(\mbox{\rm d}x).$
Now we prove that
(4.10) $\inf\limits_{z\in\mathcal{M}}\int_{\mathbb{R}^{d}}|x-z|\;\eta_{*}^{\alpha}(\mbox{\rm d}x)\leq\int_{\mathbb{R}^{d}}\operatorname{dist}(x,\mathcal{M})\,\eta_{*}^{\alpha}(\mbox{\rm d}x).$
Let $\varepsilon>0$, and let $z_{\varepsilon}$ be a quasi-minimizer for $\min\limits_{z\in\mathcal{M}}|x-z|=\operatorname{dist}(x,\mathcal{M})$, that is
$|x-z_{\varepsilon}|\leq\varepsilon+\operatorname{dist}(x,\mathcal{M}).$
Then one gets
$\int_{\mathbb{R}^{d}}|x-z_{\varepsilon}|\,\eta_{*}^{\alpha}(\mbox{\rm d}x)\leq\varepsilon+\int_{\mathbb{R}^{d}}\operatorname{dist}(x,\mathcal{M})\,\eta_{*}^{\alpha}(\mbox{\rm d}x).$
Moreover, the following always holds
$\inf\limits_{z\in\mathcal{M}}\int_{\mathbb{R}^{d}}|x-z|\,\eta_{*}^{\alpha}(\mbox{\rm d}x)\leq\int_{\mathbb{R}^{d}}|x-z_{\varepsilon}|\,\eta_{*}^{\alpha}(\mbox{\rm d}x).$
Combining the latter two inequalities, one obtains
@@
$\inf\limits_{z\in\mathcal{M}}\int_{\mathbb{R}^{d}}|x-z|\,\eta_{*}^{\alpha}(\mbox{\rm d}x)\leq\varepsilon+\int_{\mathbb{R}^{d}}\operatorname{dist}(x,\mathcal{M})\,\eta_{*}^{\alpha}(\mbox{\rm d}x).$
These quantities being independent from $\varepsilon$, we can send $\varepsilon\to 0$ and get the desired result (4.10).
Next, we continue with (4.9) which together with (4.10) yield
$\operatorname{dist}(\mathfrak{m}_{\alpha}(\rho^{\alpha}_{*}),\,\mathcal{M})$ $\leq\int_{\mathbb{R}^{d}}\operatorname{dist}(x,\mathcal{M})\,\eta_{*}^{\alpha}(\mbox{\rm d}x)$
$=\int_{\mathcal{M}_{\varepsilon}}\operatorname{dist}(x,\mathcal{M})\,\eta_{*}^{\alpha}(\mbox{\rm d}x)+\int_{\mathcal{M}_{\varepsilon}^{c}}\operatorname{dist}(x,\mathcal{M})\,\eta_{*}^{\alpha}(\mbox{\rm d}x)$
where $\varepsilon>0$ is arbitrarily fixed, and $\mathcal{M}_{\varepsilon}$ is as defined in Remark 4.6.
@@
The integrand in the first integral is upper-bounded by $\varepsilon$ since $x\in\mathcal{M}_{\varepsilon}$, which means $\operatorname{dist}(x,\mathcal{M})\leq\varepsilon$. Thus we have $\int_{\mathcal{M}_{\varepsilon}}\operatorname{dist}(x,\mathcal{M})\,\eta_{*}^{\alpha}(\mbox{\rm d}x)\leq\varepsilon$.
@@
For the second integral, we shall proceed as in the proof of the second statement in Corollary 4.4. Using Cauchy-Schwarz inequality, we have
@@
$\int_{\mathcal{M}_{\varepsilon}^{c}}\operatorname{dist}(x,\mathcal{M})\,\eta_{*}^{\alpha}(\mbox{\rm d}x)$ $\leq\bigg(\eta_{*}^{\alpha}(\mathcal{M}_{\varepsilon}^{c})\bigg)^{\frac{1}{2}}\bigg(\int_{\mathcal{M}_{\varepsilon}^{c}}\operatorname{dist}(x,\mathcal{M})^{2}\,\eta_{*}^{\alpha}(\mbox{\rm d}x)\bigg)^{\frac{1}{2}}$
@@
$\leq\bigg(\eta_{*}^{\alpha}(\mathcal{M}_{\varepsilon}^{c})\bigg)^{\frac{1}{2}}\bigg(2\,\int_{\mathcal{M}_{\varepsilon}^{c}}|x|^{2}+|z|^{2}\,\eta_{*}^{\alpha}(\mbox{\rm d}x)\bigg)^{\frac{1}{2}}\quad\text{ for some }z\in\mathcal{M}$
$\leq\bigg(\eta_{*}^{\alpha}(\mathcal{M}_{\varepsilon}^{c})\bigg)^{\frac{1}{2}}\bigg(2\,|z|^{2}+2\,\eta_{*}^{\alpha}(|\cdot|^{2})\bigg)^{\frac{1}{2}}.$
@@
$\operatorname{dist}(\mathfrak{m}_{\alpha}(\rho^{\alpha}_{*}),\,\mathcal{M})\leq\varepsilon+\bigg(\eta_{*}^{\alpha}(\mathcal{M}_{\varepsilon}^{c})\bigg)^{\frac{1}{2}}\bigg(2\,|z|^{2}+2\,\eta_{*}^{\alpha}(|\cdot|^{2})\bigg)^{\frac{1}{2}}.$
Recalling (2.1) and (3.7), the second moment $\eta_{*}^{\alpha}(|\cdot|^{2})$ is bounded as in (4.5). Finally, using Theorem 4.9 (see in particular (4.8)), we know that $\eta_{*}^{\alpha}(\mathcal{M}_{\varepsilon}^{c})\,\to 0\,$ as $\alpha\to\infty$. Here $\varepsilon$ being arbitrarily chosen, this yields $\lim\limits_{\alpha\to\infty}\operatorname{dist}(\mathfrak{m}_{\alpha}(\rho^{\alpha}_{*}),\,\mathcal{M})=0$.
Step 2. (We show $\lim\limits_{\alpha\to\infty}\,\lim\limits_{t\to\infty}\operatorname{dist}(\mathfrak{m}_{\alpha}(\rho^{\alpha}_{t}),\,\mathcal{M})=0$.)
We shall use Corollary 3.5, together with the result in Step 1. Indeed, using the triangular inequality, one has
$\operatorname{dist}(\mathfrak{m}_{\alpha}(\rho^{\alpha}_{t}),\,\mathcal{M})\leq\operatorname{dist}(\mathfrak{m}_{\alpha}(\rho^{\alpha}_{*}),\,\mathcal{M})+|\mathfrak{m}_{\alpha}(\rho^{\alpha}_{t})-\mathfrak{m}_{\alpha}(\rho^{\alpha}_{*})|.$
It suffices now to take the limit first as $t\to\infty$ using Corollary 3.5, then send $\alpha\to\infty$ using the conclusion of Step 1 above.
Step 3. (We show $\lim\limits_{\alpha\to\infty}\,\lim\limits_{t\to\infty}\operatorname{dist}\left(\kappa^{-1}\,{\mathbb{E}}\left[\,{\overline{X}}_{t}\,\right],\,\mathcal{M}\right)=0$.)
Using the triangular inequality, one has
$\operatorname{dist}\left(\kappa^{-1}\,{\mathbb{E}}\left[\,{\overline{X}}_{t}\,\right],\,\mathcal{M}\right)\leq\operatorname{dist}(\mathfrak{m}_{\alpha}(\rho^{\alpha}_{*}),\,\mathcal{M})+|\kappa^{-1}\,{\mathbb{E}}\left[\,{\overline{X}}_{t}\,\right]-\mathfrak{m}_{\alpha}(\rho^{\alpha}_{*})|.$
The conclusion thus follows using the first statement of Corollary 4.4 which holds under Assumption 2.1, together with Step 1 above. ∎